OpenStudy (blayne--moore):

Cindy has 34 meters of fencing. she plans to fence in a rectangular dog run that is 2 meters wide. [Rember, the formula for the perimeter of an rectangle (the distance around) is p=2l+2w, where L=length and w= width.] how long can the run be made?

OpenStudy (jadeishere):

 So the width of the run will not change. The length will the dependent variable (as if it's a science experiment xD) So if the width is 2 meters and there are 34 meters altogether, we can plug those two numbers in the given formula 34 = 2(2) + 2l 34 = 4 + 2l 34 - 4 = 4 -(4) + 2l 30 = 2l Divide both sides by 2 and you'll have your answer! I hope I've helped :)

OpenStudy (blayne--moore):

yeppers

OpenStudy (jadeishere):

Good! :)